Two coils A and B are connected in series across a 110V, 60Hz supply. The resistance of coil A is 2Ω and inductance of coil B is 0.004H. If the input from the supply is 2.4kW and 1.8kVAR, find the inductance of coil A and resistance of coil B. Also calculate the voltage across each coil.


1
Expert's answer
2021-03-02T09:32:03-0500

Answer





The kVA triangle and the circuit is shown above. The circuit kVA is given by, kVA="\sqrt{(2.4) ^2+(1.8) ^2}" =3

VA=3000volt -meter

Current="\frac{3000}{110}=27.27A"

So

"R_A+R_B=\frac{2400}{742.38}=" 3.22

"R_B=1.22\Omega"

Impedance of the whole circuit is given by

"Z=\frac{110}{27.27}=4.033\Omega"

So

"X_A+X_B=\sqrt{Z^2-(R_A+R_B) ^2}"


"X_A+X_B=2.42\Omega"


"X_B=2\pi \times60\times0.004=1.5\Omega"

So "X_A=0.912\Omega"

"2\pi \times60\times L_A=0.912"

"L_A=0.00241H"

So

"Z_A=\sqrt{R_A^2+X_A^2}=2.198\Omega"

P.d across coil A ="IZ_A=59.9V"


"Z_B=\sqrt{R_B^2+X_B^2}=1.933\Omega"


P.d across coil B ="IZ_B=52.72V"